What is the largest of 4 consecutive positive integers if product of the middle two integers is 2 more than 7 times the smallest integer?
Answer by stanbon(75887)

What is the largest of 4 consecutive positive integers if product of the middle two integers is 2 more than 7 times the smallest integer?
----
1st: x-2
2nd: x-1
3rd: x
4th: x+1
-------
Equation:
x(x-1) = 7(x-2)+2
----
x^2 - x = 7x - 14 + 2
-------
x^2 - 8x +12 = 0
------
(x-2)(x-6) = 0
x = 2 or x = 6
----
If x = 2 the answer is 0,1,2,3
If x = 6 the answer is 4,5,6,7
---------------
Cheers,
Stan H.
------------